OpenStudy (anonymous):

Is the line through points P(8, 3) and Q(4, 2) parallel to the line through points R(–5, –5) and S(–1, –2)? Explain. (1 point) Yes; the lines have equal slopes. Yes; the lines are both vertical. No; the lines have unequal slopes. No; one line has zero slope and the other has no slope.

OpenStudy (anonymous):

To find if lines are parallel, calculate the slope of each line. Recall that $$slope = \frac{y_2 - y_1}{x_2 - x_1} $$ Thus, $$PQslope = \frac{2-3}{4-8} = \frac{1}{4} \\ RSslope = \frac{-2 - (-5)}{-1 - (-5)} =\frac{3}{4}$$

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

Well last I checked $$\frac{1}{4} \not= \frac{3}{4}$$ :)